OpenStudy (anonymous):

solve the system using substitution 5y-x=-5 7y-x=-23

OpenStudy (anonymous):

@Nnesha @SolomonZelman

OpenStudy (anonymous):

5y-x=-5 x=5+5y 7y-(5+5y)= -23 7y-5y-5= -23 2y-5=-23 2y=-28 y=-14 x=5+5y x=5 + 5(-14) x= 5 - 70 x = -65 Proof: 5y-x=-5 5(-14) -(-65)=-5 -70+65= -5 -5 = -5 True

OpenStudy (anonymous):

thanks so much :)